(t+5)^2=4
(t+5)^2-4=0
t = -7 or -3
Please explain how you come up with this answer?

Answer by Alan3354(69443)

(t+5)^2=4
(t+5)^2-4=0
t = -7 or -3
-----------
(t+5)^2=4
t+5 = +2 --> t = -3
t+5 = -2 --> t = -7
t = -7 or -3